OpenStudy (anonymous):

Tokyo-Moto Quad Racers found that its daily profit, P, in thousands, is given by P = -4x2 + 24x - 15 where x is the number of Quad Racers manufactured daily. Which of the following is the best estimate for the maximum profit the manufacturer can earn per day, in thousands?

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

@goformit100

OpenStudy (goformit100):

@Opcode

Parth (parthkohli):

Find the derivative of the function.\[f'(x)=-8x + 24\]

Parth (parthkohli):

Equate that to zero.\[ 0 = -8x + 24\]What do you get for \(x\)?

Parth (parthkohli):

@ScienceBuddy It is obvious that you are goformit100...

Parth (parthkohli):

Also when you find the value of \(x\), find the value of \(P\) for that particular \(x\)

OpenStudy (raden):

alternative : if given an quadratic equation : ax^2 + bx + c = 0. if a < 0, the graph has the maximum value. to get the max value, u can use the formula : y = -D/4a (with D is the discriminant, defined that D = b^2-4ac) so, it can be : y = -(b^2-4ac)/4a.

OpenStudy (raden):

i mean if given f(x) = ax^2 + bx + c, it is a function not an equation